A long narrow slit of width 50  m is covered by a thin slide of glass of refractive index 1.5 and thickness of 0.5





 m, so that it covers half the slit from 0 to 25  m. Calculate the corresponding Fraunhofer diffraction pattern





when the slit is illuminated by a normally incident plane wave of wavelength 0.5  m.

1
Expert's answer
2022-01-12T08:31:28-0500

"\Delta x=\frac{d-\frac d2}{2s(n-1)}\lambda=\frac{0.5\cdot 50\cdot 10^{-6}\cdot 0.5\cdot 10^{-6}}{2\cdot 0.5\cdot 0.5\cdot 10^{-6}}=25~\mu m ."